# EMMLI: Efficient Matrix Multiplication for Machine Learning Inference through Computing in On-Chip Memory

### v2

@@ Abstract
To answer such questions regarding on-chip CiM integration for accelerating ML workloads, we use an analytical architecture-evaluation methodology where we tailor the dataflow mapping. The mapping algorithm aims to achieve highest weight reuse and reduced data movements for a given CiM prototype and workload. Next, we integrate CiM prototypes, including both analog and digital designs, into the cache levels of a tensor-core-like architecture. Our experiments show that CiM integrated memory improves energy efficiency by up to 3.4×3.4$\times$3.4 × and throughput by up to 15.6×15.6$\times$15.6 × compared to established baseline with INT-8 precision. We believe the proposed work provides insights into what type of CiM to use, and when and where to optimally integrate it in the cache hierarchy for efficient matrix multiplication.
@@ I Introduction
There are numerous ways of integrating CiM across the memory hierarchy: from CMOS on-chip cache memory to DRAM or Flash [5, 6, 7]. In this work, we focus on seamless integration of CiM in on-chip memory subsystem. While integrating CiM in caches has been explored [8, 9, 10], a comprehensive evaluation of the effectiveness of different types of CiM primitives (or designs) at the architecture level, particularly for ML inference, is yet to be studied. Our work explores the benefits of integrating CiM to different cache levels, Register File (RF) and Shared Memory (SMEM), in a Streaming Multiprocessor (SM) of a GPU [11]. To efficiently utilize CiM in a memory subsystem, there is a need to determine the optimum type of CiM, when to use it, and where to use it, for ML inference (Fig. 1). Next, we discuss the challenges in determining the answers to these questions.
@@
The rest of the paper is organized as follows: Section II distinguishes our work from other studies done in the past. Section III provides relevant background for this work. Section IV describes our dataflow-centric approach used for representing CiM enabled caches and our priority based mapping algorithm. Section V describes the experimental setup, followed by Section VI, which highlights the results and key takeaways on designing SRAM-based CiM architectures. The last section concludes the work by providing a short summary.
@@ II Related Works
While there has been work considering in-cache computing in CPUs, there have not been studies on integrating compute in GPU memory subsystem for accelerating ML inference. For example, Duality cache [8] architecture proposes last level cache in Server class Xeon processors to accelerate data parallel applications. MLIMP [9] extends the concept of Duality Cache for graph neural networks, by developing a concurrent task scheduler for multi-layer in-memory processing systems. On the other hand, this work focuses on analyzing the benefits of integrating compute at different levels of a GPU memory hierarchy for ML inference. We consider GPUs because of their widespread dominance in accelerating GEMMs, the core computation in inference tasks. In addition, GPUs are programmable accelerators and the same programming model could be potentially re-used for CiM integrated GPUs.
Livia [10] also modified cache memory in CPUs to minimize overall data movement for irregular-data-access applications. They propose a system architecture that dynamically schedules tasks and data at different locations in the memory hierarchy. In contrast, we focus on both regular and irregular workloads (GEMMs) and compare the benefits of CiM parallelism with standard architectures.
To-Pim-or-Not [21] is the first work to raise the questions on how, and when to use processing-in-memory (PIM) for different applications. It focuses on the development of a software framework to determine when and how to effectively offload computations to PIM while analyzing the trade-offs between performance benefits and offload costs. However, the scope of the work is limited to emerging general purpose DDR Memory Systems rather than SRAMs, which is the focus of our work. Another recent work [22] on benchmarking analog vs. digital compute in memory develops a quantitative energy model based on fixed analog CiM and digital CiM designs, referred to as templates. However, CiM macros vary significantly in terms of the peripheral circuits and using a template limits the design options for CiM primitive macros. Moreover, they do not provide a comprehensive evaluation on how the workload shape affects the CiM benefits.
@@ IV Dataflow-Centric Approach for representing CiM Primitives
In this section, we describe our dataflow-centric approach to represent CiM primitives, used in all subsequent evaluations. First, we provide details on how CiM is integrated in a cache memory, followed by its description in terms of CiM units. The later part of this section describes the mapping algorithm and consideration of CiM specifications in the methodology. In the end of the section, we compare our algorithm with the heuristic mapping search algorithm to highlight its benefits.
@@
Algorithm 2 Sorting Algorithm for Deciding Permutation
Three integers $m$, $n$, and $k$
if $m>n$ then
     if $n>k$ then
         return “M, N, K”
     else if $m>k$ then
         return “M, K, N”
     else
         return “K, M, N”
     end if
else
     if $n<k$ then
         return “K, N, M”
     else if $m>k$ then
         return “N, M, K”
     else
         return “N, K, M”
     end if
end if
@@
Deciding loop order: To leverage the inherent temporal reductions of CiM arrays, the loop order for compute remains as $M<K<N$, where M is in the innermost loop for highest input reuse. By changing $K$ faster than $N$, we prioritize reducing the output partial sums in the CiM primitive before moving to different partial sum. For other memory levels, priority is given to reducing the number of data accesses. Since the outermost loop affects the number of data accesses for other loops in the same memory level (Fig. 4), we follow a sorting algorithm to decide the loop order for remaining memory levels. The algorithm is explained in detail in Algo. 2. In this way, we follow a consistent dataflow mapping across different CiM primitives and GEMMs.
To incorporate the technological and circuit-level implications of CiM primitives, we follow a correct-by-construct approach by taking the costs of energy, latency and area directly from the prototypes. Since the prototypes differ in terms of supply voltage and technology, the energy numbers are scaled to match 45nm technology with 1V supply according to the established work on technology scaling [34]. The difference in frequency of operation of the CiM primitives is captured through compute cycles in terms of latency, by assuming 1GHz frequency of operation. Note that, our approach considers the essential specifications needed for deriving system level implications with CiM integrated on-chip memory. Thus, it strikes the right balance between fully accurate, rigid hardware simulation and low accuracy, flexible software modelling.
Fig. 7 shows the difference in energy efficiency ($TOPS/W$), performance ($GFLOPS$) and hardware utilization across different GEMM shapes, when compared with heuristic mapping search. The circles in the error bars represent the average change (or speedup) while the length is proportional to standard deviation in speedup. The results prove that our mapping algorithm consistently performs better in terms of utilization of resources and energy efficiency across all the GEMM shapes, including those present in ML workloads such as BERT-Large, GPT-J, ResNet50 and DLRM. As described earlier, our algorithm prioritises high utilization of CiM units by mapping largest possible weight matrix. This is in accordance with the results which show more than $\times6$ improvements in utilization, on an average, across different types of GEMMs. Given the huge mapspace, our algorithm improves energy-efficiency by 1.2×1.2$\times$1.2 × while tremendously enhancing the throughput by more than 3.2×3.2$\times$3.2 ×, on average.
@@ V-A Baseline Architecture
The baseline architecture consists of a single SM connected to main memory, akin to a GPU processing core, as shown in Fig. 1. Here, each sub-core in an SM consists of $16\times 16$ processing elements (PEs) to perform MAC operations, representing tensor-core-like operations. Further, we consider INT-8 precision, 45nm technology and 1GHz frequency of operation in all evaluations. INT-8 is chosen for its acceptable accuracy in ML inference tasks[35, 36]. The cache capacity of register file (RF) and shared memory (SMEM) are assumed to be $4\times 4~{}KB$ and $256~{}KB$ respectively. The bandwidth for SMEM and DRAM are taken as $42B$/cycle and $32B$/cycle. The energy cost of different components is displayed in Table III.
@@
V-B Workload Choices
@@ VI Results and Discussion
The purpose of this section is to address the fundamental questions posed at the beginning of the paper regarding the optimal conditions for utilizing SRAM-based CiM primitives – specifically what, when and where to integrate them. While these questions are interrelated and cannot be fully answered in isolation, we have organized this section into sub-sections focusing on each question to enhance clarity and facilitate understanding. Note, that all the results in this section adhere to the iso-area constraints, ensuring that the on-chip cache area remains the same after CiM integration. Thus, the number of CiM primitives (consequently the memory capacity) that can be accommodated in cache depends on the area overhead of the CiM primitive choice. Further, the graphs correspond to CiM integrated in register file (RF) unless stated otherwise.
@@ VI-A Effect of CiM Primitive choice
To explore the suitability of CiM primitives from system energy-efficiency and performance perspective, we choose typical designs of CiM primitives in different SRAM technology, including SRAM6T and SRAM8T. These primitives cover a range of distinct and varying parameters, as listed in the insets of Fig. 8. Each point in the graphs corresponds to a unique GEMM shape from the synthetic dataset ($GEMM(M,N,K)$ where $M,N,K$ varies from $819216~{}to~{}$ 8192). Comparing the energy efficiency, the graphs indicate that the CiM primitive with lowest energy cost ($0.09pJ$) also shows highest energy efficiency at system level, achieving more than $3~{}TOPS/W$. However, the advantage in energy efficiency (by $1-2~{}TOPS/W$) in an architecture including DRAM is significantly lower than standalone CiM accelerators which showcase two to three orders of magnitude in TOPS/W[15]. This highlights the critical bottleneck – while CiM can dramatically increase energy efficiency, the benefits are considerably constrained by the slower, more energy-intensive main memory accesses.
@@ VI-B Effect of Workload choice
Figure 9: Comparison of change in energy-efficiency, throughput and utilization for a typical digital CiM primitive integrated in register file. Each sub-graph shows how the evaluation metrics change with a particular dimension of GEMM, when other two are kept fixed at X. Particularly, (a) shows impact of weight matrix ($N\times K$) and change in $M$ dimension for each matrix size, (b) shows impact of input matrix ($M\times K$) and change in $N$ dimension, and (c) shows impact of output matrix ($M\times N$) and change in $K$ dimension for a fixed output matrix size.
@@
Figure 10: Energy-efficiency and performance for GEMMs based on different ML workloads when integrating digital CiM primitive at (a) register file (RF) and (b) shared memory (SMEM) level under iso-area constraints. configA considers an equal number of CiM primitives as the RF integration for compute. configB includes all CiM primitives that can be accommodated within the existing memory capacity of SMEM (16x compared to configA).
@@ VI-C Effect of Memory Level
Figure 11: Comparison in TOPS/W, GFLOPS and Utilization when integrating CiM at (a) RF and (b) SMEM relative to tensor-core-like baseline. The round dot shows the average change and the length corresponds to standard deviation in change across different GEMM shapes for the given ML workload. Change $>1$ implies that CiM integrated architecture performs better than baseline tensor-core-like architecture.

### v3

# What, When, Where to Compute-in-Memory for Efficient Matrix Multiplication during Machine Learning Inference

@@ Abstract
To answer such questions regarding on-chip CiM integration for accelerating ML workloads, we use an analytical architecture-evaluation methodology with tailored mapping algorithm. The mapping algorithm aims to achieve highest weight reuse and reduced data movements for a given CiM prototype and workload. Our analysis considers the integration of CiM prototypes into the cache levels of a tensor-core-like architecture, and shows that CiM integrated memory improves energy efficiency by up to 3.4×3.4$\times$3.4 × and throughput by up to 15.6×15.6$\times$15.6 × compared to established baseline with INT-8 precision. We believe the proposed work provides insights into what type of CiM to use, and when and where to optimally integrate it in the cache hierarchy for efficient matrix multiplication.
@@ I Introduction
There are numerous ways of integrating CiM across the memory hierarchy: from CMOS on-chip cache memory to DRAM or Flash [5, 6, 7]. In this work, we focus on seamless integration of CiM in on-chip memory subsystem. While integrating CiM in caches has been explored [8, 9, 10], a comprehensive evaluation of the effectiveness of different types of CiM designs at the architecture level, particularly for ML inference, is yet to be studied. Our work explores the benefits of integrating CiM to different cache levels, Register File (RF) and Shared Memory (SMEM), in a Streaming Multiprocessor (SM) of a GPU [11]. To efficiently utilize CiM in a memory subsystem, there is a need to determine the optimum type of CiM, when to use it, and where to use it, for ML inference (Fig. 1). Note, we use the term ‘primitive’ to refer to a given CiM design, inspired from the use of ‘primitives’ for linear algebra functions [12]. Next, we discuss the challenges in determining the answers to the what-when-where questions.
@@
The rest of the paper is organized as follows: Section II distinguishes our work from other studies done in the past. Section III provides relevant background for this work. Section IV describes the dataflow-centric approach used for representing CiM enabled caches and our priority based mapping algorithm. Section V describes the experimental setup, followed by Section VI, which highlights the results and key takeaways on designing SRAM-based CiM architectures. The last section concludes the work by providing a short summary.
@@ II Related Works
While in-cache computing has been explored in CPUs [8], integrating compute into the GPU memory subsystem for ML inference remains unstudied. Duality Cache [8] accelerates data-parallel applications in the last-level cache of Xeon processors, and MLIMP [9] extends this for graph neural networks with a concurrent task scheduler for multi-layer in-memory processing. On the other hand, this work focuses on analyzing the benefits of integrating compute at different levels of a GPU memory hierarchy for ML inference. We consider GPUs because of their dominance in accelerating GEMMs, the core computation in inference tasks. In addition, GPUs are programmable accelerators and the same programming model could be potentially re-used for CiM integrated GPUs.
Livia [10] also modified cache memory in CPUs to minimize data movement for irregular-data-access applications. They propose a system architecture that dynamically schedules tasks and data at different locations in the memory hierarchy. In contrast, we focus on both regular and irregular workloads (GEMMs) and compare the benefits of CiM parallelism with standard architectures.
To-Pim-or-Not [22] is the first work to raise the questions on how, and when to use processing-in-memory (PIM) for different applications, focusing on the development of a software framework for the same. However, its scope is limited to emerging general purpose DDR Memory Systems rather than SRAMs, which is the focus of our work. A recent work [23] on benchmarking analog vs. digital compute in memory develops a quantitative energy model based on fixed analog CiM and digital CiM designs, referred to as templates. However, the study lacks a comprehensive evaluation of how the workload shape affects the CiM benefits.
@@ IV Dataflow-Centric Approach for representing CiM Primitives
In this section, we describe our dataflow-centric approach to represent CiM primitives that is used in all subsequent evaluations. The term CiM primitive is used to refer to a CiM array, or in other words, an SRAM array modified to perform in-situ MAC operations. A CiM primitive further consists of multiple CiM units as explained in the first sub-section that provides details on how CiM is integrated in a cache memory. Next, we provide details on the energy and latency estimation of a CiM primitive, followed by the mapping algorithm used to schedule GEMMs on a CiM integrated architecture.
IV-A CiM representation
@@
IV-A1 CiM cost estimation
To incorporate the technological and circuit-level implications of CiM primitives, we follow an approximate approach by taking the costs of energy, latency and area directly from the prototypes. Since the prototypes differ in terms of supply voltage and technology, the energy numbers are scaled to match 45nm technology with 1V supply according to the established work on technology scaling [35]. The difference in frequency of operation of the CiM primitives is captured through compute cycles in terms of latency, by assuming 1GHz frequency of operation. The following equations describe the calculation of energy, latency and area overhead:
$Compute~{}energy~{}(pJ/mac)=\frac{2}{TOPS/W}*T_{ratio}$ (2)
$T_{ratio}=f_{45nm}/f_{ref}$ (3)
$f_{45nm}=a_{e2}^{45nm}+a_{e1}^{45nm}+a_{e0}^{45nm}$ (4)
$f_{ref}=a_{e2}*V^{2}+a_{e1}*V+a_{e0}$ (5)
[…]
$Compute~{}latency~{}(ns)=\frac{1~{}GHz}{CiM~{}frequency}*Cycles_{mac}$ (6)
$Area~{}overhead(\times)=\frac{Reference~{}CiM~{}area}{SRAM~{}area_{iso-capacity}}$ (7)
Note that, the compute energy here considers the energy of all the components comprising a CiM primitive, including the energy from the cell array, analog-digital converters (ADCs), digital-analog converters (DACs), register for bit-streaming the inputs to the array (input driver), row/col decoder, shift-and-add block, digital arithmetic logic, multiplexers, buffer for temporarily storing the MAC output and any specialized circuit block required for CiM functionality. Thus, our approach strikes the right balance between fully accurate, rigid hardware simulation and low accuracy, flexible software modelling.
@@ IV-A1 CiM cost estimation
IV-B CiM Mapping
@@
Deciding loop order: To leverage the inherent temporal reductions of CiM arrays, the loop order for compute remains as $M<K<N$, where M is in the innermost loop for highest input reuse. By changing $K$ faster than $N$, we prioritize reducing the output partial sums in the CiM primitive before moving to different partial sum. For other memory levels, priority is given to reducing the number of data accesses using a greedy algorithm. The greedy approach tries to minimize the number of data accesses locally for a memory level in the hope of reaching a globally optimal solution. Since the dimension in the outermost loop affects the number of data accesses for other dimensions in a memory level, the algorithm keeps the final loop order such that the dimension with largest loop factor is kept in the outermost for loop. For instance, in Fig. 4 (a) where $M1=3$ is in outermost dimension, all access factors are multiplied by $3$ and similarly in Fig. 4 (b) where $K1=2$ is outermost loop, all access factors have $2$ as the common factor. So, by keeping the lowest dimension in the outermost loop, we minimize the access factors locally (greedily) at each memory level.
Comparison with Heuristic Mapping: Fig. 7 shows the difference in energy efficiency ($TOPS/W$), performance ($GFLOPS$) and hardware utilization across different GEMM shapes, when compared with heuristic mapping search. The circles in the error bars represent the average change (or speedup) while the length is proportional to standard deviation in speedup. The results prove that our mapping algorithm consistently performs better in terms of utilization of resources and energy efficiency across all the GEMM shapes, including those present in ML workloads such as BERT-Large, GPT-J, ResNet50 and DLRM. As described earlier, our algorithm prioritises high utilization of CiM units by mapping largest possible weight matrix. This is in accordance with the results which show more than $\times6$ improvements in utilization, on an average, across different types of GEMMs. Given the huge mapspace, our algorithm improves energy-efficiency by 1.2×1.2$\times$1.2 × while enhancing the throughput by more than 3.2×3.2$\times$3.2 ×, on average.
@@ V-A Baseline Architecture
The baseline architecture consists of a single SM connected to main memory, akin to a GPU processing core, as shown in Fig. 1. Here, each sub-core in an SM consists of $16\times 16$ processing elements (PEs) to perform MAC operations, representing tensor-core-like operations. Note that a GPU consists of hundreds of such SMs, resulting in overall peak performance of the order of PFLOPS, while our baseline architecture consists of only 1 SM. Further, we consider INT-8 precision, 45nm technology and 1GHz frequency of operation in all evaluations. INT-8 is chosen for its acceptable accuracy in ML inference tasks[36, 37]. The cache capacity of register file (RF) and shared memory (SMEM) are assumed to be $4\times 4~{}KB$ and $256~{}KB$ respectively. The bandwidth for SMEM and DRAM are taken as $42B$/cycle and $32B$/cycle. The energy cost of different components is displayed in Table III.
@@
V-B CiM Primitive Choices
Figure 8: Examples of design choices for integrating CiM primitives in memory with (a) and (b) representing Analog-6T and Analog-8T, and (c) and (d) representing Digital-6T and Digital-8T, respectively.
TABLE IV: Single CiM primitive Specification
Sr.No. Specification
Compute Type Cell Type Rp Cp Rh Ch Capacity (KB) Latency (ns) 8b-8b MAC energy (pJ) Area (x)
1 Analog SRAM-6T 64 4 1 16 4 9 0.15 1.34
2 Analog SRAM-8T 64 4 1 16 4 144 0.09 2.1
3 Digital SRAM-6T 256 16 1 1 4 18 0.34 1.4
4 Digital SRAM-8T 1 128 10 1 4 233 0.84 1.1
The chosen primitives cover a range of distinct and varying parameters, as listed in Table IV and explained below.
The SRAM6T based analog [14] CiM primitive, illustrated in Fig. 8(a), employs local computing cells (LCCs) shared within a block of rows storing weights. LCC facilitates MAC operations between inputs passed through the input driver and the weights stored in 6T cells. Then, the analog output from LCC goes through a readout block including ADC and shift-and-add circuitry to obtain the final bits of MAC output. Here, different input bits can be fed in parallel to multiple columns, resulting in lower compute latency. However, the number of LCC blocks and ADCs limit the compute parallelism of the CiM primitive.
An example of SRAM8T based analog[15] primitive is depicted in Fig. 8(b). It can enable different number of rows (wordlines) in the array depending on the input sparsity. Such re-configurable ADC design offers high energy-efficiency, but suffers from relatively higher peripheral cost (area overhead) due to bigger ADC design. Additionally, feeding inputs in a bit-serial fashion to the same row results in a higher compute latency for such design.
The digital[16] primitive based on SRAM6T cells is shown in Fig. 8(c), that employs a fully digital design, feeding inputs into each row and executing a MAC operation at every cross-point, later combined using an adder tree. Here, the adder tree reduction incurs an area overhead and results in a higher compute latency than its analog counterpart.
One of the first digital primitives based on SRAM8T cells, shown in Fig. 8(d), displays a design where both inputs and weights are mapped to the same column. This configuration allows only two rows to be turned on at a time for a 1b-1b bitwise operation. Due to the heavily bit-serial nature of MAC computation in this design, is takes the highest compute latency but offers a minimal area overhead.
V-C Workload Choices
@@ VI Results and Discussion
The purpose of this section is to address the fundamental questions posed at the beginning of the paper regarding the optimal conditions for utilizing SRAM-based CiM primitives – specifically what, when and where to integrate them. Note, that all the results in this section adhere to the iso-area constraints, ensuring that the on-chip cache area remains the same after CiM integration. Thus, the number of CiM primitives (consequently the memory capacity) that can be accommodated in cache depends on the area overhead of the CiM primitive choice. Further, the graphs correspond to CiM integrated in register file (RF) unless stated otherwise.
@@ VI-A Effect of CiM Primitive choice
Fig. 9 compares the energy-efficiency of the chosen CiM primitives for different $GEMM(M,N,K)$ shapes from the synthetic dataset (where $M,N,K$ varies from $819216~{}to~{}$ 8192). The graphs in Fig. 9 indicate that the CiM primitive with lowest energy cost ($0.09pJ$) also shows highest energy efficiency at system level, achieving more than $3~{}TOPS/W$. However, the advantage in energy efficiency (by $1-2~{}TOPS/W$) in an architecture including DRAM is significantly lower than standalone CiM accelerators which showcase two to three orders of magnitude in TOPS/W[16]. This highlights the critical bottleneck – while CiM can dramatically increase energy efficiency, the benefits are considerably constrained by the slower, more energy-intensive main memory accesses.
@@ VI-B Effect of Workload choice
Figure 10: Comparison of change in energy-efficiency, throughput and utilization for a typical digital CiM primitive (Digital6T) integrated in register file. Each sub-graph shows how the evaluation metrics change with a particular dimension of GEMM, when other two are kept fixed at X. Particularly, (a) shows impact of weight matrix ($N\times K$) and change in $M$ dimension for each matrix size, (b) shows impact of input matrix ($M\times K$) and change in $N$ dimension, and (c) shows impact of output matrix ($M\times N$) and change in $K$ dimension for a fixed output matrix size.
@@
Figure 11: Energy-efficiency and performance for GEMMs based on different ML workloads when integrating Digital-6T CiM primitive at (a) register file (RF) and (b) shared memory (SMEM) level under iso-area constraints. configA considers an equal number of CiM primitives as the RF integration for compute. configB includes all CiM primitives that can be accommodated within the existing memory capacity of SMEM (16x compared to configA).
@@ VI-C Effect of Memory Level
Figure 12: Comparison in TOPS/W, GFLOPS and Utilization when integrating Digital-6T CiM at (a) RF and (b) SMEM relative to tensor-core-like baseline. The round dot shows the average change and the length corresponds to standard deviation in change across different GEMM shapes for the given ML workload. Change $>1$ implies that CiM integrated architecture performs better than baseline tensor-core-like architecture.
@@
TABLE V: Takeaways
Category Key observation Reason
What Maximum throughput gain is achieved by Digital-6T compared to baseline and other CiM primitives for medium to large GEMM shapes. The ability to exploit full row and column parallelism in CiM helps in improving the throughput despite higher compute latency given the GEMM shape is large enough to leverage this parallelism.
Analog-8T achieves maximum energy reduction (or highest energy efficiency) compared to baseline and other CiM architectures under is-area constraints. The overall energy consumption of CiM integrated architectures becomes proportional to their standalone MAC energy when the memory access costs are fully amortized.
When CiM integrated caches do not increase the performance of memory bound layers. The throughput of memory bound layers is limited by the main memory bandwidth.
GEMMs with high $K$ value have high energy and performance benefits from CiM integration. In contrast, small $K$ GEMMs achieve better throughput with baseline than CiM primitives. CiM primitives perform in-situ reduction over the K dimension. This reduces the number of memory accesses for partial sum reductions, but puts physical constraints on parallelism across $K$ dimension.
Where Highest performance gains are observed at SMEM level as compared to RF level under iso-area constraints. Bigger memory capacity allows for more parallelism in CiM integrated architectures and it can help surpass the cache benefits of a less parallel architecture.
The system-level energy-efficiency benefits for SMEM level are slightly higher than RF level. For large workloads that can not fully fit data dimensions in the SMEM level, the highly parallel CiM in SMEM can reduce duplicate memory accesses.